OpenStudy (anonymous):

Integrate the following tan(3x)^2 * sec(3x)^3 dx .... I'm sure there's some sort of identity I'm not remembering/seeing... I would know what to do if sec had an even power...or if tan had an odd power....

OpenStudy (anonymous):

\[tan^2x=sec^2x-1\]

OpenStudy (anonymous):

\[tan^2 3x=sec^2 3x -1\]

OpenStudy (anonymous):

\[\int\limits (\sec^5(3x) - \sec^3(3x))dx\] Is there a quick way to get to the answer without using reduction forms? Because aren't supposed to use them....

OpenStudy (anonymous):

Annnnnnnnyone know? Made it through all the other problems on my assignment, and this is the only one I'm having trouble with.

OpenStudy (perl):

derivative of tan is sec^2 , so